OpenStudy (anonymous):

A water ride at a local water park has a ride shaped like a cone that acts like a funnel whereby guests swirl around the cone until they drop through its center. There is one ride for adults and a similar, smaller version for children. If the adult ride has a radius of 32 feet and the child ride has a radius of 24 feet, what is the ratio between the volumes of each ride? 4:3 16:9 64:27 12:9

OpenStudy (ganpat):

\[1/3 \pi r ^{2} h\] - is volume of cone.. Now assuming height remains same for both of them therefore ratio turns to be.. v1/ v2 = (r1)2/ (r2)2 = 16/ 9 is the answer...